A 300v energy source delivers 500mA for 1 hour. Calculate the energy this represent.


1
Expert's answer
2021-06-20T18:13:27-0400

Gives


V=300v

I=500mA

t=1h=3600sec

Power

"P=\frac{E}{t}"

"E=P\times t"

"P=(IV)\times t"

"E=500\times 10^{-3}\times300\times3600J"

E=540000J

E=540KJ